OpenStudy (anonymous):

Anyone know an easy way to do this? If a parabola has its directrix at y = 14 and its focal point at (-4, 4), then what are the coordinates of the parabola's vertex?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How!?!?

OpenStudy (anonymous):

plug and chug into the general equation

OpenStudy (anonymous):

details ...please >.< general equation?

OpenStudy (anonymous):

(x - h)^2 = 4p(y-k) is the general equation , where p is the distance from vertex to directrix

OpenStudy (anonymous):

just sketch and you'll see

OpenStudy (anonymous):

but how do you plug in p if you don't know distance from the vertex to directrix?

OpenStudy (anonymous):

can you draw a quick sketch?

OpenStudy (anonymous):

sure :o

hero (hero):

I'll post my solution after you two are done as an alternative method.

OpenStudy (anonymous):

So whats the first step ? sorry i learn new things everyday. I'm barely learning this today so bare with me..

OpenStudy (anonymous):

 draw the porabola for me

OpenStudy (anonymous):

draw it ._. how!?

OpenStudy (anonymous):

click on "Draw" button, then click "copy previous drawing" , then draw

OpenStudy (anonymous):

i tried but it's blank?

OpenStudy (anonymous):

?!?!

hero (hero):

Okay, I'm just going to post my solution now. Give me a minute.

OpenStudy (anonymous):

@sourwing :o

hero (hero):

Okay, so basically, you are given two points \((-4, 4)\) and \((x, 14)\). Plug them in to this formula: \((x - x_1)^2 + (y - y_1)^2 = ((x - x_2)^2 + (y - y_2)^2\) \((x - (-4))^2 + (y - 4)^2 = (x - x)^2 + (y - 14)^2\) \((x + 4)^2 + (y - 4)^2 = (y - 14)^2\) Expand \((y - 4)^2\) and \((y - 14)^2\) \((x + 4)^2 + y^2 - 8y + 16 = y^2 - 28y + 196\) \(y^2\) cancels on both sides leaving just: \((x + 4)^2 - 8y + 16 = y^2 - 28y + 196\) Add 8y to both sides; Subtract 196 from both sides: \((x + 4)^2 + 16 - 196 = -28y + 8y\) \((x + 4)^2 - 180 = -20y\) Divide both sides by -20 \(-\dfrac{(x + 4)^2}{20} + 9 = y\) or y = \(-\dfrac{1}{20}(x + 4)^2 + 9\) Notice that the parabola is of the form \(y = a(x - h)^2 + k\) where \((h, k)\) is the vertex. Thus, the coordinates of the vertex of the parabola is \((-4, 9)\).

OpenStudy (anonymous):

Okay this looks a little easier for me.. I barely learned how to graph parabolas today so let me try...

hero (hero):

That's because it's a completely algebraic method. No graphing necessary.

OpenStudy (anonymous):

thanks for your time^.^ im gonna try it.. so starting off, is it (x-(-4)^2 + (y-4)^2 = ....

hero (hero):

Yes, (x - (-4))^2 + (y - 4)^2 = ... keep going

OpenStudy (anonymous):

yay ok.....

OpenStudy (anonymous):

ok but see.. this is where i get stuck.. what is (x-x2)^2 supposed to be?

hero (hero):

Notice that the second point which represents the directix is \((x_2, y_2) = (x , 14)\). For \((x - x_2)^2\) You replace \(x_2\) with \(x\) to get \((x - x)^2\)

OpenStudy (anonymous):

okay got it. so , so far so good right ? just keep going.. (x-x2)^2 then becomes (x-x)^2 ?

hero (hero):

You're probably wondering why we're writing the directrix as a point instead of a line. It's because the point \((x , 14)\) is the same as the line y = 14 because no matter what the value of x, y will always be 14.

OpenStudy (anonymous):

nope i got it.... kinda weird for a second but then i figured it was the same as y=14 lol

hero (hero):

Which makes it very convenient to use the distance formula for this.

OpenStudy (anonymous):

yes!! im glad you gave me this formula

hero (hero):

You can use it no matter what the direction of the parabola is.

hero (hero):

I made a slight mistake in my steps above. I said \(y^2\) cancels on both sides at a certain point, which it does, but I cancelled on one side and forgot to cancel on the other side. Please ignore that typo. The rest of it is corect as written.

OpenStudy (anonymous):

ok! ... so i got (x+4)^2+(y-4)^2=(x-x)^2+(y-14)^2 ... correct so far?

hero (hero):

Yes correct. Notice that \((x - x)^2 = (0)^2 = 0\) since \(x - x = 0\)

OpenStudy (anonymous):

oh thanks....i was just gonna say how do you solve that...

OpenStudy (anonymous):

I got (2x+16)+(2y+16)=(2y+196) ?

OpenStudy (anonymous):

Omg you wrote the same thing ._. i didn't know you were actually writing the same problem

hero (hero):

I don't know how you got that. There's no way to get that from your previous step.

OpenStudy (anonymous):

I'm confusing myself

OpenStudy (anonymous):

I hate math i give up :(

hero (hero):

\((y - 4)^2 = (y - 4)(y - 4) \) \(\\= y(y - 4) - 4(y - 4) \\= y^2 - 4y - 4y + 16 \\= y^2 - 8y + 16\)

hero (hero):

Never give up

OpenStudy (anonymous):

Now what >.<

hero (hero):

Just review the steps I posted (minus the typo)

OpenStudy (anonymous):

I'm going to study your steps allllll night

OpenStudy (anonymous):

thanks so much for thoroughly explaining. i appreciate it soooooo much!

hero (hero):

You're welcome